# Hilbert schemes on blowing ups and the free Boson

### v2

@@ Abstract.
For different cohomology theories (including the Hochschild homology, Hodge cohomology, and Chow groups), we identify the cohomology of moduli space of rank $1$ perverse coherent sheaves on the blow-up of a surface by Nakajima-Yoshioka [22] as the tensor product of cohomology of Hilbert schemes with a Fermionic Fock space. As the stable limit, we identify the cohomology of Hilbert schemes of blow-up of a surface as the tensor product of cohomology of Hilbert schemes on the surface with a Bosonic Fock space. The actions of the infinite-dimensional Clifford algebra and Heisenberg algebra are all given by geometric correspondences.
@@ 1.1. Motivation and the Main Theorem
Let $S$ be a smooth projective surface over $\mathbb{C}$. The Poincaré polynomials of Hilbert schemes of points on $S$ were computed by the famous Göttsche’s formula:
@@
    By replacing the cohomology with the Chow group, the Heisenberg (and moreover, Virasoro and $W_{1+\infty}$) algebra action was constructed by Maulik-Negut [19]. However, as the Chow group of a surface can be infinite-dimensional (like $K3$ surfaces), the Chow groups of all $S^{n}$ are invovled to decide the Chow group of Hilbert schemes, and the representation theory is also more complicated (for example, $sp_{2\infty}$ will naturally appear).
@@
    where $B$ is the bosonic Fock space $\mathbb{Q}[y_{1},y_{2},\cdots]$. The formula (1.2) does not depend on the choice of the surface $S$ and the point $o$, and might also hold if we replace the cohomology group with Chow groups and Hochschild homology.
@@
Let $\mathbb{H}^{*}$ be the functor of Hodge cohomology groups, Hochschild homology groups, or the Chow groups. We have an identification
(1.3) $\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}B$
@@
More details of our geometric correspondences will be given from Section 1.2 to Section 1.4. At the current stage, we do not know the precise relation between our Heisenberg algebra action with that of Nakajima [21] and Grojnowski [10] in cohomlogy groups or Maulik-Neguţ [19] in Chow groups.
@@ 1.2. Stable limits and perverse coherent sheaves
The right way to consider Theorem 1.1 is to regard the Bosonic Fock space $B$ and $\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})$ as a stable limit:
@@
    We consider the infinite dimensional vector space $V$ and the Fermionic Fock space
    $V=\bigoplus_{i=1}^{\infty}\mathbb{Q}x_{i},\quad F:=\bigoplus_{i=0}^{\infty}\wedge^{i}V.$
    There is a natural grading on
    $F=\bigoplus_{i=0}^{\infty}F_{i}$
    by associating every $x_{i}$ to degree $1$. We consider the degree $1$ operator $K_{1}$ maps
    $x_{j_{1}}\wedge\cdots\wedge x_{j_{k}}\to x_{0}\wedge x_{j_{1}+1}\wedge\cdots\wedge x_{j_{k}+1};$
    Then $K_{1}$ is injective and induces a stable limit $F_{\infty}:=\lim_{i\to\infty}F_{i}$. A careful study of $F_{\infty}$ will reveal that
    $F_{\infty}\cong B.$
@@
    Let $p:\hat{S}\to S$ be the projection. Nakajima-Yoshioka [22] considered the moduli space of perverse coherent sheaves on $\hat{S}$, which connects $S^{[n]}$ and $\hat{S}^{[n]}$ by the following properties:
    1. (a)
      fixing a chern character in $H^{*}(\hat{S},\mathbb{Q})$
      $c_{n}=[\hat{S}]+n[pts]$
      there exists moduli space $M^{m}(c_{n})$ of perverse coherent sheaves for any integer $m\geq 0$ such that
      $M^{0}(c_{n})\cong S^{[n]},\quad M^{r}(c_{n})\cong\hat{S}^{[n]}\text{ if }r\gg 0.$
    2. (b)
      $M^{r}(c_{n})$ and $M^{r+1}(c_{n})$ are birational and connected through a wall-crossing.
    Thus $\coprod_{n=0}^{\infty}\hat{S}^{[n]}$ can be regarded as a stable limit of
    $\coprod_{n=0}^{\infty}M^{m}(c_{n}).$
Theorem 1.1 follows from the geometric representation theory of perverse coherent sheaves:
@@
(1.4) $\bigoplus_{m=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{m}(c_{n}),\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}F$
and the action of the infinite dimensional Clifford algebra on $F$ is given by geometric correspondences.
@@ 1.3. A categorical Boson-Fermion correspondence
The fact that the bosonic Fock space $B$ is the stable limit of the Fermionic Fock space $F$ is a special case of a categorical Boson-Fermion correspondence which is of independent interest. It reveals a equivalence of certain infinite dimensional Heisenberg algebra (which we denote as $\mathcal{H}$) representations and graded infinte dimensional Clifford algebra (which we denote as $\mathcal{C}$) representations. The bridge between $\mathcal{C}$ and $\mathcal{H}$ is an $\mathbb{Q}$-algebra, which we call $\mathcal{E}$:
@@
Then $P_{i},Q_{i}$ satisfy the relations of infinite dimensional Clifford algebras (see Proposition 2.2 for the details). On the other hand, the Fermionic Fock space $F$ can also be regarded as the highest weight representation of $\mathcal{E}$ generated by a vector $|0\rangle$ and the relations (see Proposition 2.8 for the details)
$F|0\rangle=0,\quad K_{-1}|0\rangle=K_{1}|0\rangle=|0\rangle.$
@@ 1.4. Derived Grassmannians and incidence varieties
By the categorical Boson-Fermion correspondence, what we would like to construct is an $\mathcal{E}$-module structure on
@@
by incidence varieties. It follows from the incidence variety of the derived Grassmannian which is introduced by Jiang [11]: given a locally free sheaf $\mathcal{I}$ over a variety $X$, the Grassmannian (under the notation of Grothendieck) $\mathcal{I}$, which we denotes as $Gr(\mathcal{I},r)$, is the moduli space of quotients
@@
where $\mathcal{E}$ is a rank $r$ locally free sheaf on $X$. Jiang [11] generalized the concept of the Grassmannian and allows that $\mathcal{I}$ to be a perfect complex, i.e. locally a complex of locally free sheaves. The derived Grassmannian has good properties when $\mathcal{I}$ has tor-amplitude $[0,1]$, i.e. it is locally a two-term complex of locally free sheaves. Particularly, in this situation $\mathcal{I}^{\vee}[1]$ is also a tor-amplitude complex. For example, if $\mathcal{I}$ is the two term complex $V\to W$, then $\mathcal{I}^{\vee}[1]$ is $W^{\vee}\to V^{\vee}$.
@@
Hence the indicence variety of derived Grassmannians induce the maps
@@ 1.5. Some related work
Although not explicitly pointing it our paper, the representation we constructed is closed related to the quiver Yangian action of Li-Yamazaki [18] and Galakhov-Li-Yamazaki [8]. Here we take an example that $S$ is the equivariant $\mathbb{C}^{2}$ (although it is not projective). Li-Yamazaki [18] defined the affine super-Yangian of $gl(1|1)$, which we denote as
@@
When $S$ is the equivariant $\mathbb{C}^{2}$, the moduli space of perverse coherent sheaves was described as a quiver-like variety by Nakajima-Yoshioka [23]. The affine super-Yangian $Y_{\hbar_{1},\hbar{2}}(\widehat{gl(1|1)})$ acts on the cohomology of moduli spaces of perverse coherent sheaves by Galakhov-Morozov-Tselousov [9]. The operators $\phi_{l-1,l},\phi_{l,l},\psi_{l-1,l},\psi_{l,l}$ in Theorem 1.6 actually comes from the action of certain elements in $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$, which we will discuss in the future work.
@@ 1.5.2. Cohomological Hall algebra (COHA)
By the dimension reduction, the affine super-Yangian is also the double COHA of certain one-dimensional objects over a surface, which was defined by Diaconescu-Porta-Sala [6]. A systematically computation of the double COHA will appear very soon [7].
@@ 1.5.3. Koseki’s categorification
Another work which we get great inspiration is the categorification of the derived category of coherent sheaves on the Hilbert scheme of points [15] by Koseki. To get more precise representation theoretic relations, we apply the semi-orthogonal decomposition theorem of Jiang [13] with explicit Fourier-Mukai kernels.
@@ 1.6. Notations in this paper
All the varieties and derived schemes in this paper will be over $\mathbb{C}$.
@@ 1.7. Acknowledgements
I would like to thank Hiraku Nakajima, Andrei Neguţ, Francesco Sala, Qingyuan Jiang, Jiakang Bao, Dmitry Galakhov and Henry Liu for many interesting discussions on the subject.
@@ 2. A categorical Boson-Fermion correspondence
Consider the infinite-dimensional Clifford algebra $\mathcal{C}$ generated by $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{>0}}$ satisfying the anti-commutator relations
(2.1) $\{P_{i},P_{j}\}=0,\quad\{Q_{i},Q_{j}\}=0,\quad\{P_{i},Q_{j}\}=\delta_{ij}$
and the infinite-dimensional Heisenberg algebra $\mathcal{H}$ generated by $\{H_{i}\}_{i\in\mathbb{Z}-\{0\}}$ with the commutator relations
$[H_{i},H_{j}]=i\delta_{i,-j}$
In this section, we demonstrate an equivalence between specific representations of $\mathcal{C}$ and $\mathcal{H}$ which can be regarded as a categorical version of a Boson-Fermion correspondence.
The equivalence is established through a new algebra, denoted as $\mathcal{E}$, generated by $K_{1},K_{-1},E,F$ with the relations
@@
The algebra $\mathcal{E}$ naturally emerges from the (semi)-orthogonal decomposition of vector spaces or linear categories: given a decomposition of a vector space
$V\cong V_{0}\oplus V_{1},$
let $F$ (resp. $E$) be the projection (resp. embedding) from $V$ to $V_{1}$ (resp. from $V_{1}$ to $V$). Let $K_{-1}$ (resp. $K_{1}$) be the projection (resp. embedding) from $V$ to $V_{0}$ (resp. $V_{0}$ to $V$). The operators satisfy the relations 2.2.
@@ 2.1. Basic properties of $\mathcal{E}$ and $\mathcal{E}$-representations
The elements $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{\geq 0}}$ satisfy the relations of $\mathcal{C}$, i.e. 2.1.
@@
Given a $\mathcal{E}$-module $W$, the action of $K_{1},E$ on $W$ are injective and the action of $K_{-1},F$ on $W$ are surjective. Moreover,
$K_{1}W\cong ker(F),\quad EW\cong ker(K_{-1}),\quad W\cong K_{1}W\oplus EW.$
as vector spaces.
@@
The injective (resp. surjective) property of $K_{1},E$ (resp. $K_{-1},F$) follows from
$K_{-1}K_{1}=FE=1.$
Now we prove that $K_{1}W\cong ker(F)$ (and similarly $EW\cong Ker(K_{-1})$). First we have $K_{1}W\subset ker(F)$ as $FK_{1}=0$. Second, for any $f\in Ker(F)$, we have
$f=K_{1}K_{-1}f+EFf=K_{1}(K_{-1}f)\in K_{1}W.$
Moreover, if $K_{-1}f=0$, then $f=0$ and hence
$ker(f_{0})\cap ker(f_{1})=K_{1}W\cap EW=0$
and thus we have $W\cong K_{1}W\bigoplus EW$. ∎
@@
Any non-zero $\mathcal{E}$-module has to be infinite-dimensional: for a finite-dimensional representation, we have
$K_{1}K_{-1}=id,\quad EF=id,$
which contradicts with that $K_{1}K_{-1}+EF=id$.

### v3

@@ Abstract.
For different cohomology theories (including Hochschild homology, Hodge cohomology, Chow groups, and Grothendieck groups of coherent sheaves), we identify the cohomology of the moduli space of rank $1$ perverse coherent sheaves on the blow-up of a surface by Nakajima-Yoshioka [22] as the tensor product of the cohomology of Hilbert schemes with a Fermionic Fock space. As the stable limit, we identify the cohomology of Hilbert schemes of the blow-up of a surface as the tensor product of the cohomology of Hilbert schemes on the surface with a Bosonic Fock space. The actions of the infinite-dimensional Clifford algebra and Heisenberg algebra are all given by geometric correspondences.
@@ 1.1. Motivation and the Main Theorem
Let $S$ be a smooth projective surface over $\mathbb{C}$. The famous Göttsche’s formula computed the Poincaré polynomials of Hilbert schemes of points on $S$ [11]:
@@
    By replacing the cohomology with the Chow group, the Heisenberg (and moreover, Virasoro and $W_{1+\infty}$) algebra action was constructed by Maulik-Negut [20]. However, as the Chow group of a surface can be infinite-dimensional (like $K3$ surfaces), the Chow groups of all $S^{n}$ are involved to determine the Chow group of Hilbert schemes, and the representation theory is also more complicated (for example, $sp_{2\infty}$ will naturally appear).
@@
    where $B$ is the Bosonic Fock space $\mathbb{Q}[y_{1},y_{2},\cdots]$. The formula (1.2) does not depend on the choice of the surface $S$ and the point $o$, and might also hold if we consider the generalized cohomology theory like Hochschild homology, Chow groups and Grothendieck group of coherent sheaves.
@@
Let $\mathbb{H}^{*}$ be the functor of Hodge cohomology groups, Hochschild homology groups, Chow groups, or Grothendieck groups of coherent sheaves. We have an identification
(1.3) $\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}B,$
@@
More details of our geometric correspondences will be given in Section 1.2 to Section 1.4. At the current stage, we do not know the precise relation between our Heisenberg algebra action and that of Nakajima [21] and Grojnowski [10] in cohomology groups or Maulik-Neguţ [20] in Chow groups. We expect the study of the affine super-Yangian will shed some light on this problem. See Section 1.5 for more discussions.
@@ 1.2. Stable Limits and Perverse Coherent Sheaves
Theorem 1.1 follows from regarding both the Bosonic Fock space $B$ and $\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})$ as a stable limit:
@@
    We consider the Fermionic Fock space $F$ which is the exterior algebra of
    $V=\bigoplus_{i=1}^{\infty}\mathbb{Q}x_{i}.$
    The degree $1$ operator $K_{1}$
    $K_{1}:F_{i}\to F_{i+1},\quad x_{j_{1}}\wedge\cdots\wedge x_{j_{i}}\to x_{0}\wedge x_{j_{1}+1}\wedge\cdots\wedge x_{j_{i}+1}$
    is injective. Moreover, by direct computation, the stable limit
    $F_{\infty}:=\lim_{i\to\infty}F_{i}\cong B.$
@@
    Nakajima-Yoshioka [22] considered the moduli space of $l$-stable perverse coherent sheaves $M^{l}(c_{n})$ on $\hat{S}$, where $l\in\mathbb{Z}$ and
    $c_{n}=[\hat{S}]+n[o]$
    is the Chern character. It has the property that
    $M^{0}(c_{n})\cong S^{[n]},\quad M^{l}(c_{n})\cong\hat{S}^{[n]}\text{ if }l\gg 0.$
    Hence $\hat{S}^{[n]}$ can be regarded as the stable limit of $M^{l}(c_{n})$, $l\to\infty$.
Theorem 1.1 follows from the representation theory on $M^{l}(c_{n})$:
@@
(1.4) $\bigoplus_{l=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}F$
and the action of the infinite-dimensional Clifford algebra on $F$ is given by geometric correspondences.
@@ 1.3. A categorical Boson-Fermion correspondence
The fact that the Bosonic Fock space $B$ is the stable limit of the Fermionic Fock space $F$ is a special case of a categorical Boson-Fermion correspondence which is of independent interest. It reveals an equivalence of certain infinite-dimensional Heisenberg algebra (which we denote as $\mathcal{H}$) representations and graded infinite-dimensional Clifford algebra (which we denote as $\mathcal{C}$) representations. The bridge between $\mathcal{C}$ and $\mathcal{H}$ is an $\mathbb{Q}$-algebra, which we call $\mathcal{E}$:
@@
Then $P_{i},Q_{i}$ satisfy the relations of infinite-dimensional Clifford algebras (see Proposition 2.2 for the details). On the other hand, the Fermionic Fock space $F$ can also be regarded as the highest weight representation of $\mathcal{E}$ generated by a vector $|1\rangle$ and the relations (see Proposition 2.7 for the details)
$F|1\rangle=0,\quad K_{-1}|1\rangle=K_{1}|1\rangle=|1\rangle.$
@@ 1.4. Derived Grassmannians and incidence varieties
Due to the categorical Boson-Fermion correspondence, we would like to construct an $\mathcal{E}$-module structure on
@@
by incidence varieties. It follows from the incidence variety of the derived Grassmannian which is introduced by Jiang [12]: given a locally free sheaf $\mathcal{I}$ over a variety $X$, the Grassmannian (under the notation of Grothendieck) $\mathcal{I}$, which we denote as $Gr(\mathcal{I},r)$, is the moduli space of quotients
@@
where $\mathcal{E}$ is a rank $r$ locally free sheaf on $X$. Jiang [12] generalized the concept of the Grassmannian and allows $\mathcal{I}$ to be a perfect complex, i.e., locally a complex of locally free sheaves. A special case we shall care about is when $\mathcal{I}$ has tor-amplitude $[0,1]$, i.e., it is locally a two-term complex of locally free sheaves. Particularly, in this situation $\mathcal{I}^{\vee}[1]$ is also a tor-amplitude complex.
@@
Hence the indicence variety of derived Grassmannians induces the maps
@@ 1.5. Some related work
Although not explicitly pointed out in our paper, the representation we constructed is closely related to the quiver Yangian action of Li-Yamazaki [19] and Galakhov-Li-Yamazaki [8]. Here we take an example where $S$ is the equivariant $\mathbb{C}^{2}$ (although it is not projective). Li-Yamazaki [19] defined the affine super-Yangian of $gl(1|1)$, which we denote as
@@
When $S$ is the equivariant $\mathbb{C}^{2}$, the moduli space of perverse coherent sheaves was described as a quiver-like variety by Nakajima-Yoshioka [23]. The affine super-Yangian $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$ acts on the cohomology of moduli spaces of perverse coherent sheaves by Galakhov-Morozov-Tselousov [9]. The operators in Theorem 1.6 actually come from the action of certain elements in $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$, which we will discuss in future work.
@@ 1.5.2. Cohomological Hall algebra (COHA)
By dimension reduction, the affine super-Yangian is also the double COHA of certain one-dimensional objects over a surface, which was defined by Diaconescu-Porta-Sala [6]. A systematic computation of the double COHA will appear very soon [7].
@@ 1.5.3. Koseki’s categorification
Another work from which we draw great inspiration is the categorification of the derived category of coherent sheaves on the Hilbert scheme of points by Koseki [16]. To obtain more precise representation-theoretic relations, we apply the semi-orthogonal decomposition theorem of Jiang [14] with explicit Fourier-Mukai kernels.
@@ 1.6. Notations in this paper
1.6.1. Derived algebraic geometry
All the varieties and derived schemes in this paper will be over $\mathbb{C}$. For a derived scheme $X$, we will denote $Perf(X)$ as the $\infty$-category of perfect complexes on $X$ and $QCoh(X)$ as the $\infty$-category of quasi-coherent complexes.
1.6.2. Representation theory
@@
We denote $\mathcal{C}$ as the infinite-dimensional Clifford algebra, i.e., the $\mathbb{Q}$-algebra generated by $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{>0}}$ with relations
(1.6) $\{P_{i},P_{j}\}=0,\quad\{Q_{i},Q_{j}\}=0,\quad\{P_{i},Q_{j}\}=\delta_{ij}.$
We also denote $\mathcal{H}$ as the infinite-dimensional Heisenberg algebra, i.e., the $\mathbb{Q}$-algebra generated by $\{H_{i}\}_{i\in\mathbb{Z}-\{0\}}$ with relations
$[H_{i},H_{j}]=i\delta_{i,-j}.$
We denote $B$ as the Bosonic Fermion space, i.e. the polynomial ring of infinitely many variables $\mathbb{Q}[y_{1},y_{2},\cdots]$. It is generated by variables $y_{i},i\in\mathbb{Z}_{>0}$ with the relation
$[y_{i},y_{j}]=0,\forall i,j>0.$
Let $U:=\oplus_{i\in\mathbb{Z}_{>0}}\mathbb{Q}y_{i}$, then $B$ is the symmetric algebra of $U$. We also denote $F$ as the Fermionic Fock space, i.e. the super-polynomial ring of infinitely many variables $\mathbb{Q}[x_{1},x_{2},\cdots]$. It is generated by variables $x_{i},i\in\mathbb{Z}_{>0}$ with the relation
$\{x_{i},x_{j}\}=0,\forall i,j>0.$
Let $V:=\oplus_{i\in\mathbb{Z}_{>0}}\mathbb{Q}x_{i}$, then $F$ is the exterior algebra of $V$.
@@ 1.7. Acknowledgements
I want to thank Hiraku Nakajima, Andrei Neguţ, Francesco Sala, Qingyuan Jiang, Jiakang Bao, Dmitry Galakhov, and Henry Liu for many interesting discussions on the subject.
@@ 2. A categorical Boson-Fermion correspondence
In this section, we demonstrate an equivalence between specific representations of $\mathcal{C}$ and $\mathcal{H}$ which can be regarded as a categorical version of a Boson-Fermion correspondence. The equivalence is established through a new algebra, denoted as $\mathcal{E}$, generated by $K_{1},K_{-1},E,F$ with the relations
@@
It naturally emerges from the orthogonal decomposition of a vector space
$V\cong V_{0}\oplus V_{1}.$
Then the respective operators of projections and embedding
$F:V\to V_{1},\quad E:V_{1}\to V,\quad K_{-1}:V\to V_{0},\quad K_{1}:V_{0}\to V$
satisfy the relations 2.1.
@@ 2.1. Basic properties of $\mathcal{E}$ and $\mathcal{E}$-representations
The elements $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{\geq 0}}$ satisfy the relations of $\mathcal{C}$, i.e. 1.6.
@@
Given an $\mathcal{E}$-module $W$, the actions of $K_{1}$ and $E$ on $W$ are injective, and the actions of $K_{-1}$ and $F$ on $W$ are surjective. Moreover, as vector spaces,
$K_{1}W=\ker(F),\quad EW=\ker(K_{-1}),\quad W=K_{1}W\oplus EW.$
@@
The respective injective and surjective properties follow from
(2.3) $K_{-1}K_{1}=FE=1.$
As $K_{-1}E=FK_{1}=0$, we have
$K_{1}W\subset\ker(F),\quad EW\subset\ker(K_{-1}).$
As $K_{1}K_{-1}+EF=1$, we have
$\ker(F)=(K_{1}K_{-1}+EF)\ker(F)=K_{1}K_{-1}\ker(F)\subset K_{1}W.$
Hence $\ker(F)=K_{1}W$ and similarly $EW=\ker(K_{-1})$. Still by the fact that $K_{1}K_{-1}+EF=1$, we have
$W=(K_{1}K_{-1}+EF)W=K_{1}K_{-1}W+EFW\subset K_{1}W+EW,$
$\ker(F)\cap\ker(K_{-1})\subset\ker(K_{1}K_{-1}+EF)=0$
and thus $W=K_{1}W\oplus EW$. ∎
@@
Any non-zero $\mathcal{E}$-module is infinite-dimensional: as $K_{1}$ and $W$ are injective, we have
$W\cong K_{1}W\oplus EW\cong W\oplus W.$